how do I do this problem: 4x+5y+6x
Found 2 solutions by checkley71, elima:
Answer by checkley71(8403)

4X+5Y+6X=10X+5Y=5(2X+Y) ANSWER.

Answer by elima(1433)

4x+5y+6x
I'm not sure what your directions are as far as solving. Are you suppose to solve for x, for y or just collect like terms?
The way it is right now it looks like you are suppose to just collect like terms;
4x + 5y + 6x
10x + 5y